OpenStudy (anonymous):

2. What is the range of the relation? {(1, 2); (2, 4); (3, 1), (4, 3)} (Points : 4) {1, 2, 3, 4} {1, 2, 3, 4, 6} {2, 4, 6} {2, 4}

OpenStudy (anonymous):

set of y co-ordinates, in this case: {1,2,3,4}

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

3. For the function f(x) = 2 − 3x, find f(4). (Points : 4) −10 −5 12 14

OpenStudy (anonymous):

just put 4 in place of x

OpenStudy (anonymous):

2-3*4=2-12=-10

OpenStudy (anonymous):

i was just going to say -10

OpenStudy (anonymous):

4. Write the equation of the function graphed. VHS_ALG_L3_S2_01_16_UTQ4.gif (Points : 4) f(x) = |x| + 4 f(x) = |x| − 4 f(x) = |x + 4| f(x) = |x − 4|

OpenStudy (anonymous):

there is no link. please post separately for separate questions. cheers :)

OpenStudy (anonymous):

VHS_ALG_L3_S2_01_16_UTQ4.gif